# Proxy-Free GFlowNet

### v1

@@ Appendix B Related works
GFlowNets [1, 2] are a class of generative models designed to sample compositional objects from an unnormalized reward distribution via sequential action selection. They have been theoretically linked to entropy-regularized reinforcement learning [24, 45] and variational inference [46, 47], offering a principled framework for distribution matching. Recent research has focused on improving GFlowNet training by proposing new balance conditions [42, 48], enhancing credit assignment through intermediate rewards [37, 49, 9], designing backward policies for improved sampling [34, 45, 32, 50], refining sampling and resampling strategies [23, 10, 51, 52, 53], and exploring alternative training objectives [54, 55]. Extensions to the GFlowNet framework have expanded its applicability to stochastic dynamics [56], continuous action spaces [57], non-acyclic transitions [58], and implicit reward feedback [59]. These advances have broadened the scope of GFlowNets, enabling impactful applications in molecule design [18, 60, 61, 62], biological sequence generation [3, 63], combinatorial optimization [4, 64, 65], causal inference [66, 67], recommendation systems [68], and fine-tuning of large language models [5, 69] and diffusion models [6, 70].
@@
Inverse reinforcement learning (IRL) aims to recover a reward function that explains observed expert behavior in a Markov decision process [71]. Among its variants, maximum causal entropy IRL is particularly noteworthy, as it addresses reward ambiguity by formulating the objective as finding a policy that maximizes entropy while matching the observed expert behavior [25, 26]. Building on this framework, Generative Adversarial Imitation Learning (GAIL) [28] reinterprets imitation as a distribution matching problem via adversarial training, where a discriminator distinguishes expert from policy-generated trajectories, and its output serves as a learned reward for policy training. GAIL eliminates the need to explicitly solve the inner IRL loop, significantly improving scalability, and has inspired a broad class of Adversarial Imitation Learning (AIL) methods [72, 29, 73, 74]. This line of work has since been extended in various directions, including efforts to improve reward generalization and robustness [75, 76].
@@ C.1 Hypergrid
The Hypergrid task is a synthetic benchmark introduced in [1] to evaluate the exploration and generalization capabilities of GFlowNet algorithms. It presents a significant challenge due to its high-dimensional, sparse reward landscape.
@@
Following the most challenging configuration proposed in [1], we set the environment parameters to $R_{0}=10^{-3},R_{1}=0.5,R_{2}=2$. This configuration results in a highly multi-modal and sparsely rewarded environment, making it well-suited for evaluating the effectiveness of offline training strategies.
@@ C.2 Biosequence Design
The Anti-Microbial Peptide (AMP) Design task, introduced in [3], is a realistic sequence generation benchmark aimed at designing short peptide sequences with strong anti-microbial properties.
The agent constructs sequences by selecting amino acids from a fixed vocabulary of $20$ standard residues, with a maximum sequence length of $50$. At each time step, the agent chooses either an amino acid or a special end-of-sequence token, proceeding in a left-to-right manner. As a result, the transition structure forms a tree.
@@
In our experiments, we adopt the Diversity metric to evaluate algorithmic performance, following the methodology in [3]. For a set of biological sequences $C$, the metric is defined as:
@@
where $\text{Lev}(\cdot,\cdot)$ denotes the Levenshtein distance. All environmental parameters are kept consistent with those used in [3].
@@ C.3 Molecule Design
We evaluate our method on a fragment-based molecular generation task targeting protein binding affinity, specifically for soluble epoxide hydrolase (sEH). The goal is to construct candidate molecules with high predicted affinity for the target protein. Molecules are generated sequentially using the junction tree framework introduced in [77], where each action involves selecting an attachment atom and adding a fragment from a predefined vocabulary of $105$ building blocks.
The reward is defined as the negative binding energy between the generated molecule and the target protein, as predicted by an oracle model pre-trained on $300k$ molecules, following the setup in [1]. To balance reward maximization with output diversity, the final reward is modulated using a reward exponent $\omega$, such that $R^{\prime}(x)=R(x)^{\omega}$, where $R(x)$ denotes the raw oracle prediction. All environmental parameters, including $\omega$, are kept consistent with those used in [1].
@@ C.4 Policy Model Architectures and Hyperparameters
Molecule Design MPNN [78] v4
Key algorithmic hyperparameters used across different tasks and methods are summarized in Table 3. These values were selected through empirical tuning to ensure stable training and fair comparisons among baselines. [^1] Our implementation of COFlowNet is based on the official open-source repository available at https://github.com/yuxuan9982/COflownet. Specifically, the pruning threshold coefficient $K$ is tuned by selecting the value that yields the highest ratio of retained expert transitions to retained randomly selected transitions.
@@ D.1 Algorithm trained on extremely limited data
To further assess the robustness of TD-GFN under extreme data scarcity, we conduct experiments on the $H=8$, $D=4$ Hypergrid environment using only $30$ training trajectories, as illustrated in Figure 7. In this setting, algorithms are prone to overfitting due to the limited data. For example, both COFlowNet and CQL exhibit non-monotonic trends in Empirical L1 Error, initially decreasing but later increasing during training, indicating overfitting to the limited known regions.
@@
D.2 Algorithm Visualization on Hypergrid
@@
D.3 Time consumption and resource usage
@@ D.4 Ablation Study
D.5 Orthogonality to Training Paradigms
As shown in Algorithm 1, the improvements introduced by TD-GFN are orthogonal to the specific training paradigm employed after trajectory collection. While we adopt the Flow-Matching (FM) objective [1] in this work, our framework is also compatible with alternative GFlowNet training objectives—such as DB [2], TB [42], and SubTB [48]—without requiring substantial modification.
Due to time and space constraints, we focus our empirical analysis on the integration of TD-GFN with offline RL methods and the quantile matching technique proposed in [38].
Replacing Flow Matching with Offline RL Training
@@ D.5 Orthogonality to Training Paradigms
Instead of using the standard flow-matching objective in GFlowNet training, we examine the effect of replacing it with the Conservative Q-Learning (CQL) objective from offline reinforcement learning [35], resulting in a variant we refer to as TD-CQL. We evaluate TD-CQL on both the Expert and Mixed datasets, each consisting of $1,500$ trajectories. Results are presented in Figure 11.
@@
With training components inherited from TD-GFN, TD-CQL successfully discovers all modes using a comparable number of state visits and achieves an Empirical L1 Error below $10^{-4}$ with fewer visits than TD-GFN, demonstrating strong sample efficiency and early-stage accuracy in modeling the target distribution. However, beyond this point, the Empirical L1 Error increases, indicating a decline in distributional approximation quality. This degradation arises from the nature of the offline RL objective, which tends to overfit the policy toward only the highest-reward objects, thereby limiting its ability to accurately model the full reward distribution.
@@
Figure 11: Performance comparison of TD-CQL on the Expert and Mixed datasets, each consisting of $1,500$ trajectories.
Incorporating Quantile Matching
@@
Following the approach of [22], we augment TD-GFN with the quantile matching technique proposed in [38], resulting in a variant referred to as QM-TD-GFN. We evaluate QM-TD-GFN on the Molecule Design task (Section 4.3) to assess its impact on the diversity of generated samples.
@@
D.6 Discussion on Baseline Performance
@@
Although CQL is also an offline RL algorithm, its performance differs from IQL due to a distinct training paradigm. Unlike IQL, which learns exclusively from transitions in the dataset, CQL explicitly constrains policy actions to regions supported by the dataset by penalizing overestimation of Q-values. As a result, its performance is highly sensitive to dataset quality. A similar sensitivity is observed in COFlowNet, which applies conservative regularization to restrict the policy’s behavior to the data distribution’s support.
@@
Appendix E Future Directions
While TD-GFN demonstrates strong empirical performance and improved robustness over existing GFlowNet training methods, several limitations remain, suggesting promising directions for future research.
First, TD-GFN relies on an adversarial approach to estimate edge-level rewards, which can introduce training instability and requires careful hyperparameter tuning. Developing more stable and generalizable techniques for edge reward estimation—potentially drawing on recent advances in inverse reinforcement learning—could enhance the reliability and scalability of the framework.
Second, the current formulation of TD-GFN assumes a deterministic environment. Extending the method to accommodate stochastic rewards or non-deterministic transitions [56] is a critical direction, particularly for real-world applications where uncertainty is inherent in both reward evaluation and environment dynamics.

### v2

@@ Appendix B Related work
GFlowNets (Bengio et al., 2021; 2023) are a class of generative models designed to sample compositional objects from an unnormalized reward distribution via sequential action selection. They have been theoretically linked to entropy-regularized reinforcement learning (Tiapkin et al., 2024; Mohammadpour et al., 2024) and variational inference (Malkin et al., 2023; Zimmermann et al., 2023), offering a principled framework for distribution matching.
Recent research has focused on improving GFlowNet training by proposing new balance conditions (Malkin et al., 2022; Madan et al., 2023), enhancing credit assignment through intermediate rewards (Pan et al., 2022; 2023a; Jang et al., 2023), designing backward policies for improved sampling (Jang et al., 2024; Mohammadpour et al., 2024; Shen et al., 2023; Gritsaev et al., 2024), refining sampling and resampling strategies (Atanackovic & Bengio, 2024; Kim et al., 2024c; b; Lau et al., 2024; Madan et al., 2025), and exploring alternative training objectives (Silva et al., 2024; Hu et al., 2024b), including energy-based training approaches for diffusion-structured inference (Campbell et al., 2024). Extensions to the GFlowNet framework have expanded its applicability to stochastic dynamics (Pan et al., 2023b), continuous action spaces (Lahlou et al., 2023), non-acyclic transitions (Brunswic et al., 2024; Morozov et al., 2025), and implicit reward feedback (Chen & Mauch, 2024). These advances have broadened the scope of GFlowNets, enabling impactful applications in molecule design (Jain et al., 2023a; b; Zhu et al., 2023; Roy et al., 2023), biological sequence generation (Jain et al., 2022; Ghari et al., 2023), combinatorial optimization (Zhang et al., 2025a; 2023a; Kim et al., 2024a), causal inference (Zhang et al., 2022; Deleu et al., 2022), recommendation systems (Liu et al., 2023), and fine-tuning of large language models (Hu et al., 2024a; Yu et al., 2024) and diffusion models (Liu et al., 2025; Venkatraman et al., 2024).
Our work is most closely related to recent efforts in offline and proxy-free GFlowNet training. An early approach, RO-GFlowNets (Wang et al., 2023), uses offline action probabilities to constrain the learned policy, while COFlowNet (Zhang et al., 2025b) penalizes flow on unseen edges to limit policy coverage to dataset-adjacent regions. Both methods aim to improve policy reliability by imposing coarse constraints that merely align policy behavior or coverage more closely with the dataset. However, such constraints impose blunt limitations that hinder generalization and make performance highly sensitive to data quality, as they fail to fully exploit the trajectory-level information contained in the dataset. Other related works include Atanackovic & Bengio (2024), which explores an “offline” setting but primarily focuses on evaluating sampling strategies, and the KL-weakFM loss (Brunswic et al., 2025), which enables proxy-free imitation learning. Concurrently, some research has investigated improving offline training with explicit reward signals and novel exploration strategies (Sendera et al., 2024), providing complementary perspectives to our pruning-based framework.
Conceptually, our work shares its core insight with Silva et al. (2025), which introduces Weighted Detailed Balance (WDB) to emphasize transitions with a larger impact on correctness, particularly in online GFlowNets with GNN-based parameterizations. Their analysis demonstrates that imbalances across edges can systematically affect sampling accuracy. Our work shares this motivation but takes a different route: instead of reweighting balance conditions, we employ inverse reinforcement learning (IRL) to estimate edge rewards. These rewards are then used to guide pruning and prioritized sampling in the offline setting. This distinction in methodology makes the two approaches complementary—WDB could potentially enhance offline training, while our pruning and sampling strategies may in turn benefit online methods.
@@
Inverse reinforcement learning (IRL) aims to recover a reward function that explains observed expert behavior in a Markov decision process (Ng et al., 2000). Among its variants, maximum causal entropy IRL is particularly noteworthy, as it addresses reward ambiguity by formulating the objective as finding a policy that maximizes entropy while matching the observed expert behavior (Ziebart et al., 2008; 2010). Building on this framework, Generative Adversarial Imitation Learning (GAIL) (Ho & Ermon, 2016) reinterprets imitation as a distribution matching problem via adversarial training, where a discriminator distinguishes expert from policy-generated trajectories, and its output serves as a learned reward for policy training. GAIL eliminates the need to explicitly solve the inner IRL loop, significantly improving scalability, and has inspired a broad class of Adversarial Imitation Learning (AIL) methods (Fu et al., 2018; Kostrikov et al., 2018; 2020; Garg et al., 2021). This line of work has since been extended in various directions, including efforts to improve reward generalization and robustness (Xu et al., 2020; Luo et al., 2022; 2024).
@@ C.1 Hypergrid
The Hypergrid task is a synthetic benchmark introduced in Bengio et al. (2021) to evaluate the exploration and generalization capabilities of GFlowNet algorithms. It presents a significant challenge due to its high-dimensional, sparse reward landscape.
@@
Following the most challenging configuration proposed in Bengio et al. (2021), we set the environment parameters to $R_{0}=10^{-3},R_{1}=0.5,R_{2}=2$. This configuration results in a highly multi-modal and sparsely rewarded environment, making it well-suited for evaluating the effectiveness of offline training strategies.
@@ C.2 Biosequence Design
The Anti-Microbial Peptide (AMP) Design task, introduced in Jain et al. (2022), is a realistic sequence generation benchmark aimed at designing short peptide sequences with strong anti-microbial properties.
The agent constructs sequences by selecting amino acids from a fixed vocabulary of $20$ standard residues, with a maximum sequence length of $50$. At each time step, the agent chooses either an amino acid or a special end-of-sequence token, proceeding in a left-to-right manner. As a result, the transition structure forms a tree, and the trajectories in offline datasets can be directly inferred by the terminal states.
@@
In our experiments, we adopt the Diversity metric to evaluate algorithmic performance, following the methodology in Jain et al. (2022). For a set of biological sequences $C$, the metric is defined as:
@@
where $\text{Lev}(\cdot,\cdot)$ denotes the Levenshtein distance. All environmental parameters are kept consistent with those used in Jain et al. (2022).
@@ C.3 Molecule Design
We evaluate our method on a fragment-based molecular generation task targeting protein binding affinity, specifically for soluble epoxide hydrolase (sEH). The goal is to construct candidate molecules with high predicted affinity for the target protein. Molecules are generated sequentially using the junction tree framework introduced in Jin et al. (2018), where each action involves selecting an attachment atom and adding a fragment from a predefined vocabulary of $105$ building blocks.
The reward is defined as the negative binding energy between the generated molecule and the target protein, as predicted by an oracle model pre-trained on $300k$ molecules, following the setup in Bengio et al. (2021). To balance reward maximization with output diversity, the final reward is modulated using a reward exponent $\omega$, such that $R^{\prime}(x)=R(x)^{\omega}$, where $R(x)$ denotes the raw oracle prediction. All environmental parameters, including $\omega$, are kept consistent with those used in Bengio et al. (2021).
@@ C.4 Policy Model Architectures and Hyperparameters
Molecule Design MPNN (Gilmer et al., 2017) v4
@@
Key algorithmic hyperparameters used across different tasks and methods are summarized in Table 3. These values were selected through empirical tuning to ensure stable training and fair comparisons among baselines. [^1] Our implementation of COFlowNet is based on the official open-source repository available at https://github.com/yuxuan9982/COflownet. Specifically, the pruning threshold coefficient $K$ is tuned by selecting the value that yields the highest ratio of retained recorded transitions to retained randomly selected transitions. For the sensitivity analysis with respect to parameter $K$, please refer to Appendix D.4.
@@ D.1 Algorithm trained on extremely limited data
In fact, our experiments in 4.1 have explicitly tested and demonstrated robustness of TD-GFN under sufficiently sparse or unevenly distributed data. Specifically, the sampled $150$ trajectories from the Mixed dataset and $1,500$ trajectories from the Bad dataset cover only $\mathbf{12}$ and $\mathbf{6}$ out of $16$ reward modes, respectively. To further assess the robustness of TD-GFN under extreme data scarcity, we conduct experiments on the $H=8$, $D=4$ Hypergrid environment using only $\mathbf{30}$ training trajectories from the Expert/Mixed datasets which cover merely $\mathbf{5}$ and $\mathbf{3}$ modes respectively, as illustrated in Figure 8. In this setting, algorithms are prone to overfitting due to the limited data. For example, both COFlowNet and CQL exhibit non-monotonic trends in Empirical L1 Error, initially decreasing but later increasing during training, indicating overfitting to the limited known regions.
@@
D.2 Algorithm trained on a larger Hypergrid
We also evaluate the performance of TD-GFN in a more challenging setting with sparser rewards, using a larger $20^{4}$ Hypergrid environment. As shown in Figure 9, TD-GFN achieves strong performance across all four datasets, consistently and significantly outperforming COFlowNet.
Figure 9: Performance comparison on the $20^{4}$ Hypergrid task using $1,500$ trajectories from four types of datasets.
D.3 Algorithm Visualization on Hypergrid
@@
D.4 Sensitivity analysis of the pruning threshold coefficient $K$
In our method, the pruning threshold hyperparameter $K$ is used to guide edge removal in a statistically principled manner (Equation (6)). It is tuned by selecting the value that yields the highest ratio of retained recorded transitions to retained randomly selected transitions in this work. Our experiments demonstrate that applying a fixed K(=7)K\,(=7) across different datasets within the same environment yields strong performance without the need for dataset-specific tuning. While future work may explore more theoretically grounded strategies for DAG pruning, here we provide a sensitivity analysis of this parameter. As shown in Figure 11, our method remains robust across different values of $K$ without catastrophic degradation.
Figure 11: Performance comparison between different values of the pruning coefficient $K$ on the $8^{4}$ Hypergrid task. Policies are trained using four types of datasets, each consisting of $1,500$ trajectories.
D.5 Time consumption and resource usage
@@ D.6 Ablation Study
D.7 Orthogonality to Training Paradigms
As shown in Algorithm 1, the improvements introduced by TD-GFN are orthogonal to the specific training paradigm employed after trajectory collection. Due to time and space constraints, our empirical analysis primarily focuses on the compatibility of TD-GFN with existing GFlowNet training objectives, as well as its integration with offline RL methods and the quantile matching technique proposed by Zhang et al. (2023b).
Compatibility with GFlowNet training objectives.
With trajectories sampled on the pruned DAG $G^{\prime}=(V^{\prime},E^{\prime})$, we can readily apply various GFlowNet training objectives to optimize our policy on this subgraph without requiring substantial modifications, such as the flow matching (FM) objective (Bengio et al., 2021) and the trajectory balance (TB) objective (Malkin et al., 2022), as illustrated below.
By setting $R(s)=0$ for all non-terminal states $s$, the FM objective for each sampled trajectory $\tau$ is given by:
The TB objective for each sampled trajectory $\tau$ is given by:
$\mathcal{L}_{TB}(\tau;\theta)=\left(Z_{\theta}+\sum_{(s\rightarrow s^{\prime})\in E^{\prime}}\big(\mathcal{P}_{F}^{G^{\prime}}(s^{\prime}\mid s;\theta)-\mathcal{P}_{B}^{G^{\prime}}(s\mid s^{\prime};\theta)\big)-R(x)\right)^{2},$ (9)
where $\mathcal{P}_{F}^{G^{\prime}}(s^{\prime}\mid s;\theta)$ and $\mathcal{P}_{B}^{G^{\prime}}(s\mid s^{\prime};\theta)$ denote the forward and backward transition probabilities respectively, both normalized over the edges in the pruned subgraph $G^{\prime}$ is the reward of the terminal state $x$, and $Z_{\theta}$ is the learned normalizing constant. Moreover, we find it beneficial to fix $\mathcal{P}_{B}^{G^{\prime}}$ to the form given in Equation (7), a practice also noted in Malkin et al. (2022). The results of applying the TB objective on the four datasets used in our main experiments are presented in Figure 14.
Figure 14: Performance comparison of TD-GFN with different training objectives as well as COFlowNet on the $8^{4}$ Hypergrid task. Policies are trained using four types of datasets, each consisting of $1,500$ trajectories.
We observe that the TB objective converges faster than the FM objective, but is less effective in rapidly exploring modes and exhibits weaker stability against overfitting on extreme datasets (e.g., Bad). Nevertheless, it still substantially outperforms COFlowNet.
Replacing GFlowNet objectives with Offline RL Training.
@@ D.7 Orthogonality to Training Paradigms
Instead of using the standard objectives in GFlowNet training, we examine the effect of replacing it with the Conservative Q-Learning (CQL) objective from offline reinforcement learning (Kumar et al., 2020), resulting in a variant we refer to as TD-CQL. We evaluate TD-CQL on both the Expert and Mixed datasets, each consisting of $1,500$ trajectories. Results are presented in Figure 15.
@@
With training components inherited from TD-GFN, TD-CQL successfully discovers all modes using a considerably smaller number of state visits compared to CQL and achieves an Empirical L1 Error below $10^{-4}$ with fewer visits than TD-GFN, demonstrating strong sample efficiency and early-stage accuracy in modeling the target distribution. However, beyond this point, the Empirical L1 Error increases, indicating a decline in distributional approximation quality. This degradation arises from the nature of the offline RL objective, which tends to overfit the policy toward only the highest-reward objects, thereby limiting its ability to accurately model the full reward distribution.
@@
Figure 15: Performance of GFlowNet and CQL with and without our Trajectory-Distilled (TD) framework on the Expert and Mixed datasets, each consisting of $1,500$ trajectories.
Incorporating Quantile Matching.
@@
Following the approach of Zhang et al. (2025b), we augment TD-GFN with the quantile matching technique proposed in Zhang et al. (2023b), resulting in a variant referred to as QM-TD-GFN. We evaluate QM-TD-GFN on the Molecule Design task (Section 4.3) to assess its impact on the diversity of generated samples.
@@
D.8 Discussion on Baseline Performance
@@
Although CQL is also an offline RL algorithm, its performance differs from IQL due to a distinct training paradigm. Unlike IQL, which learns exclusively from transitions in the dataset, CQL explicitly constrains policy actions to remain close to the data by penalizing the overestimation of Q-values. As a result, its performance is highly sensitive to dataset quality. A similar sensitivity is observed in COFlowNet, which applies conservative regularization to limit its policy coverage to dataset-adjacent regions.
@@
Appendix E Discussions
Why Can TD-GFN Surpass an Oracle-GFN?
It is indeed surprising that TD-GFN achieves superior performance compared to a GFlowNet trained directly with an oracle reward model in the Molecule Design task. We offer two primary hypotheses to explain this outcome.
The first hypothesis relates to the exploration efficiency of the learning frameworks. A proxy-based GFlowNet, including one guided by an oracle, learns via a trial-and-error process of exploration and subsequent policy updates. Given the vastness of the search space, it is plausible that this exploration only covers a fraction of the high-reward areas that the oracle is aware of, leading the final policy to concentrate on a limited set of modes. TD-GFN, by distilling guidance from offline trajectories, may learn a more comprehensive exploration strategy that generalizes better beyond the specific samples in the dataset to a wider range of high-reward candidates.
The second hypothesis concerns the manner in which reward signals are utilized. The process of training a proxy or oracle aims to approximate the reward function over the entire space, treating different reward values with a uniform objective (e.g., minimizing mean squared error). In contrast, TD-GFN’s methodology is more aligned with backward planning, which naturally prioritizes valuable goals. By rebalancing the dataset and using prioritized backward sampling, our framework explicitly emphasizes high-reward terminal states. The resulting edge rewards learned through IRL are not a simple reward surrogate but a form of dense, structural guidance, potentially highlighting more efficient paths to top-tier candidates within that chemical space.
Adaptability to Extended GFlowNet Settings.
TD-GFN can be directly applied to scenarios involving non-acyclic directed graphs (Brunswic et al., 2024), as it operates over general transition structures without a fundamental reliance on acyclicity. In contrast, adapting our framework to continuous environments would necessitate additional modeling considerations (e.g., for defining and pruning over continuous edge sets). Furthermore, extending the method to handle stochastic rewards or non-deterministic transitions (Pan et al., 2023b) represents an important direction, particularly for real-world applications where uncertainty is inherent in both reward evaluation and environmental dynamics.
Toward More Advanced Edge Reward Modeling.
The current implementation of TD-GFN employs an adversarial IRL approach for edge reward estimation. This choice, while effective, inherits the well-known challenges of adversarial training, such as potential instability and sensitivity to hyperparameters. Therefore, the development of more stable IRL techniques presents a promising avenue for further enhancing the reliability and scalability of the TD-GFN framework. This may be achieved by leveraging recent advances in inverse reinforcement learning, such as stability-enhancing regularization (Fu et al., 2018; Kostrikov et al., 2020), representation learning–augmented IRL (Chandak et al., 2019), and score-based or energy-based IRL methods (Liu et al., 2020).
On Dataset Generation.
Many of the datasets in our work are generated using GFlowNets. This strategy is also common in the offline reinforcement learning (RL) community (Fu et al., 2020; Gulcehre et al., 2020), where datasets are typically collected by RL policies of varying quality. However, our method is not tied to any particular data generation procedure. For instance, in Section 4.2, we employ a dataset curated from the DBAASP database (Pirtskhalava et al., 2021). We adopt GFlowNet-generated datasets not because our approach requires them, but because they provide a controlled and reproducible way to simulate realistic scenarios in which data arises from behavior policies of different optimality levels. This allows for a systematic investigation of how dataset quality influences performance. Moreover, certain real-world datasets were not included simply because no proxy-based learning baselines are available for comparison on them. In future work, we aim to further validate the effectiveness of our approach on datasets that more closely resemble practical deployment environments.
What is Encoded in Trajectory Data.
Our work is situated in the offline GFlowNet setting, where training is conducted on a fixed collection of reward-labeled trajectories. These trajectories—whether collected from humans, heuristics, or simulators—are often overlooked in proxy-based methods. In contrast, our approach is designed to leverage such information without resorting to proxy reward models.
In this setting, however, our method differs from proxy-based approaches in that it explicitly leverages the trajectories themselves and assumes that they encode valuable information beyond terminal rewards. A natural question, then, is what constitutes a valuable trajectory. To deepen our understanding of the role of trajectories in offline GFlowNet training, we construct two additional datasets from the terminal states used in the main experiments by generating trajectories in reverse: (i) Uniform, where at each step a parent node is selected uniformly at random; and (ii) Rule-based, where the parent is chosen by preferentially reducing the dimension with the smallest current value, thereby producing trajectories that avoid the central regions of the hypergrid with low rewards. We then apply TD-GFN to these two datasets and compare the results with those obtained using the original GFlowNet-generated dataset. The results are presented in Figure 18.
Figure 18: Performance comparison of TD-GFN policies trained on the $8^{4}$ Hypergrid task using three types of datasets with the same terminal states, each consisting of $1,500$ trajectories.
Although all three datasets share the same set of terminal states, the results clearly show that the trajectories themselves provide crucial information. Trajectories generated via uniform backward sampling perform very poorly, and even the rule-based trajectories, which are highly efficient and deliberately avoid low-reward regions, are less effective than trajectories generated forward by a GFlowNet. This observation suggests TD-GFN leverages meaningful trajectory structure, implying it may struggle with datasets consisting only of noise. Nevertheless, its sensitivity is mild compared to COFlowNet, which degrades more significantly under such conditions.
Future Directions
The observation above opens up two promising directions for future work. First, developing principled methods for generating effective trajectories from terminal states could extend TD-GFN’s applicability to settings where only terminal data is available. Second, the strong performance of TD-GFN on GFlowNet-generated datasets suggests its potential in online training, where it could dynamically prune the DAG using the agent’s own experience to improve efficiency for continued training. Furthermore, an interesting future direction would be to replace the “hard” pruning of the DAG with a “softer” guidance mechanism, such as re-weighting the policy logits using the learned edge rewards.